# Precise quantum-geometric electronic properties from first principles

### v1

@@ 3.2 Gallium arsenide
The GaAs bands near $\Gamma$ illustrate how the degenerate perturbation theory described in the previous section works in practice. In the $[111]$ direction ($\Gamma-$L), diagonalization of the $4\times 4$ matrix $A^{(1)}(\hat{\bf q}=[111],n=\text{VBM},{\bf k}=\Gamma)$ of Eq. (21) for the four-fold degenerate valence band maximum (VBM) gives two nonzero and nondegenerate band velocities with opposite signs, and a twice-degenerate vanishing band velocity (see left inset of Fig. 2). For the two nondegenerate band velocities, the masses are obtained from the relevant diagonal elements of $\tilde{A}^{(2)}([111],\text{VBM},\Gamma)$, which results from applying to $A^{(2)}([111],\text{VBM},\Gamma)$ the unitary transformation that diagonalizes $A^{(1)}([111],\text{VBM},\Gamma)$; see Eqs. (33) to (36). These two diagonal elements are identical, implying that the corresponding masses are identical. For the doubly-degenerate level with vanishing band velocity, the masses are obtained by diagonalizing the corresponding $2\times 2$ submatrix of $\tilde{A}^{(2)}([111],\text{VBM},\Gamma)$, yielding two identical masses. In the $[100]$ direction ($\Gamma-$X), $A^{(1)}([100],\text{VBM},\Gamma)$ has two doubly-degenerate nonzero band velocities with opposite signs (see middle inset of Fig. 2). For each sign, the corresponding $2\times 2$ submatrix of $\tilde{A}^{(2)}([100],\text{VBM},\Gamma)$ must be diagonalized; the final result is four identical masses as discussed previously in connection with Fig 3. Finally, in the $[110]$ direction ($\Gamma-$K), $A^{(1)}([110],\text{VBM},\Gamma)$ has four distinct band velocities: two with large absolute values, and two with small absolute values (see right inset of Fig. 2). All the masses are obtained from the corresponding diagonal elements of $\tilde{A}^{(2)}([110],\text{VBM},\Gamma)$. The fact that at $\Gamma$ the effective mass degeneracies are multiples of two is a consequence of time reversal symmetry.
@@ 4 Gapped graphene
Gapped graphene is a system where the inversion symmetry of the two-dimensional pristine structure is broken by an external potential, typically associated with a substrate on which graphene is grown epitaxially [44, 45]. The equivalence between the two atomic sites in the primitive unit cell is broken, and as a result a gap opens in the Dirac cones at the Fermi level. The simplest tight-binding (TB) model [46, 47], with just one $p_{\text{z}}$ orbital per site, leads to a two-band low-energy model with just two parameters: the magnitude of the gap $\Delta$ and the Fermi velocity $v_{\text{F}}=\sqrt{3}at/2$ [46], where $a$ is the lattice constant and $t$ the hopping integral of the TB model. Quantum-geometric quantities have simple analytical expressions in two-band models [48, 49], and the low-energy model for gapped graphene has been widely used in recent times to investigate such quantities [50].
@@
The self-consistent potential of broken-symmetry graphene was calculated using a modified Becke-Johnson meta-GGA functional [43] in a supercell configuration. The Tran-Blaha constant [43] was fixed at $c_{\text{TB}}=1.04$, since the original recipe is inadequate for isolated 2D materials. While the LDA approximation is known to underestimate the bonding energy of $\sigma$ electrons with respect to $\pi$ electrons [51], we find that the modified Becke-Johnson meta-GGA functional gives results closer to experiment. A Troullier-Martins [38] pseudopotential was constructed from a local-density ground-state configuration, with a core radius of 1.3 atomic units for the $s$, $p$, and $d$ channels. The local potential was a smoothed maximum of all channels, and the nonlocal part was converted to the Kleynman-Bylander form. The self-consistent calculation used a kinetic energy cutoff of 50 Ha for the plane-wave expansion, and a $6\times 6\times 2$ uniform grid for BZ integration. The lattice constant in the calculations was $2.456\,\,\text{\AA}$ in the graphene plane, three times that value was used in the perpendicular supercell direction. The added inversion-breaking potential was tweaked to open a gap of 0.28 eV; this is the same value that was used in Ref. [50], and is also very close to the experimental value reported for epitaxial graphene on a SiC substrate [44].
@@
Figure 4 shows the DFT-pseudopotential energy bands as a function of wavevector distance $q$ from K towards $\Gamma$, together with the bands of the low-energy model,
$E_{\pm}=\pm\frac{\Delta}{2}\sqrt{1+(q/q_{0})^{2}},$ (38)
@@
$\Omega=\pm\frac{1}{2q_{0}^{2}}\frac{1}{\left[1+(q/q_{0})^{2}\right]^{3/2}};$ (40)
@@
with $x$ in the direction from K to $\Gamma$. At K $g^{xx}=g^{yy}=|\Omega|/2$ [note the factor-of-two difference in Eqs. (23) and (24)], and the results from the model with the fitted $q_{0}$ are very close to the DFT values, particularly for $q_{x}\ll q_{0}$. As in the case of the Berry curvature, the differences in the DFT values for the valence and conduction bands are so small that they are not noticeable on the scale of the figure. The dependence on $q_{x}$ is very different for the longitudinal and transverse components of the quantum metric. The model fit of the DFT results is very good for the longitudinal component, but for the transverse component there is a noticeable deviation. As the system has a third dimension, the $zz$ component of the metric has a nonzero value, $g^{zz}=0.56\,\,\text{\AA}^{2}$ at $q=0$, which is very small on the scale of Fig. 5, and depends very weakly on $q_{x}$.
@@
$\Bigl{(}\frac{1}{m^{*}}\Bigr{)}^{xx}(q_{x},0)$ $=\frac{1}{m_{e}}\frac{\Delta}{2q_{0}^{2}}\frac{1}{\left[1+(q_{x}/q_{0})^{2}\right]^{1/2}}$ (42)
$\Bigl{(}\frac{1}{m^{*}}\Bigr{)}^{yy}(q_{x},0)$ $=\frac{1}{m_{e}}\frac{\Delta}{2q_{0}^{2}}\frac{1}{\left[1+(q_{x}/q_{0})^{2}\right]^{3/2}},$
with $m_{e}$ the electron mass (one in atomic units). The small deviations seen in Fig. 4 between the DFT and model energy bands get magnified in the inverse masses, which are their second derivatives. For $q_{x}=0$, there is a small difference in masses between the valence and conduction band that is not apparent on the scale of the figure. As the average effective mass was used for the value of $q_{0}$ in the DFT fit, the agreement between the model and DFT calculations is not surprising. When moving away from K towards $\Gamma$, the longitudinal and transverse inverse masses become different, the longitudinal mass decreasing at a much faster rate.
@@
where the Bohr magneton, $\mu_{\text{B}}$, has the value $1/2$ in atomic units. In contrast with previous figures, there is a clear difference between the orbital magnetic moments of the valence and conduction bands at K, although they both have values (in atomic units) close to half the inverse mass, as predicted by the low-energy model. The sign of the orbital magnetic moment is the same for the two bands, but, like the Berry curvature, it depends on which of the two inequivalent K points is considered, or on which atom the symmetry-breaking potential is placed. A consistency check is that the orbital moment must have the same sign as the Berry curvature of the conduction band.
@@
Figure 7: The relations between the trace and determinant of the quantum metric $g$ and the magnitude of the Berry curvature $\Omega$ are illustrated by plotting differences between those quantities as a function of the distance from K towards $\Gamma$. Notice the change of scale with respect to Fig. 5. In the top panel, the dashed-black lines and solid-green lines both pertain to the low-energy model and are obtained by inserting slightly different sets of parameters into Eq. (41). For the quantity plotted in the bottom panel, that model predicts a null value.
@@
Next, we analyze the close connection between the real and imaginary parts of the quantum-geometric tensor, i.e., between the quantum metric and the Berry curvature. For nondegenerate states and two-dimensional materials, the quantum-geometric tensor becomes a $2\times 2$ hermitian matrix. The three invariants of such a matrix are the determinant and trace of the real-symmetric part (quantum metric), and the value of the imaginary-antisymmetric part (Berry curvature). From the positive semi-definiteness of the quantum-geometric tensor, one obtains the following inequalities between the three invariants [52, 53],
$\frac{1}{2}\operatorname{tr}(g)\geq\sqrt{\det{g}}\geq\frac{1}{2}|\Omega|.$ (44)
@@
(The first inequality is just between the arithmetic and geometric averages of the non-negative eigenvalues of the quantum metric.) Inspection of Eqs. 40 and 41 shows that the second inequality saturates for the low-energy model (this actually happens for any two-band model [53]), while the first only saturates in the limits $q\ll q_{0}$ and $q\gg q_{0}$. In Fig. 7, the quantities $\frac{1}{2}\operatorname{tr}(g)-\sqrt{\det{g}}$ and $\sqrt{\det{g}}-\frac{1}{2}|\Omega|$ are plotted as a function of wavevector distance from K towards $\Gamma$. Comparing the vertical scale of this figure with that of Fig. 5 shows that the inequalities are very close to being saturated, particularly in the case of $\sqrt{\det{g}}-\frac{1}{2}|\Omega|$ where the two-band model would predict a null value.
@@
The two-band low-energy model fails to take into account the presence of other bands in the DFT calculation, which are responsible for the small nonzero values displayed in the bottom panel of Fig. 7. To estimate the corrections to the two-band model, note that the sum-over-states expression for $Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\alpha}\rangle$ in Eq. 14 contains an energy-difference denominator. The Berry curvature and quantum metric will therefore depend on the inverse of the square of those energy differences. The smallest among them is the gap $\Delta=0.28\,\,\text{eV}$; the next relevant level (taking mirror symmetry into account) is about 9 eV away, and therefore $9/0.28\sim 32$ times farther away. The corrections to the metric and Berry curvature should therefore be of the order of $(1/32)^{2}$, or about 0.1%. Indeed, this is the factor between the vertical scale of the bottom of Fig. 7 and that of Fig. 5. In the case of the inverse mass and orbital moment, the sum-over-states expressions depend on the inverse of the energy differences (not squared); as such, deviations from the two-band model are expected to be more pronounced, of the order of 3%. That is in fact almost the value by which the orbital magnetic moments at K in Fig. 6 deviate from their average. The difference between the absolute values of the conduction and valence effective masses is 1.2%.
@@ 5 Trigonal tellurium
[…]
@@
Figure 8: The energy bands of Te near the band gap are shown along the H–K line, near the H point ($q_{z}=0$). The upper valence and lower conduction bands are shown in the left and right panels, respectively. On this highly detailed scale, the “camel back” structure of the valence band is clearly seen. The inset shows the hexagonal Brillouin zone. The short cyan segment on the H–K line highlights the considered region.
@@
The valence-band maximum (VBM) of trigonal Te, relevant for the naturally $p$-doped material, is located on the H–K line, the lateral edge of the hexagonal BZ, very close to the H points, which form the vertices of the hexagonal BZ: see inset of Fig. 8. The band dispersion displays a “camel back” feature which is clearly seen in the left panel of Fig. 8, with a local minimum at H (a saddle point in three dimensions), and a maximum about $0.02\,\,\text{\AA}^{-1}$ close to it. In the present calculation, the depth of the minimum with respect to the maximum is 1.3 meV. The experimental value is 1.1 meV [57], the calculated value with an LAPW code and also a modified Becke-Johnson functional is 1.7 meV [16], and the calculated value with the VASP code and the HSE06 functional is 0.8 meV [18]. These are very small energies, and while the order of magnitude is consistent, there is some variation associated with different computational methods and functionals.
@@
Figure 9: The Berry curvature (top panel) and quantum metric (bottom panel) of the upper valence band of Te are shown along the H–K line as a function of distance from the H point. In the top panel the solid line is the Berry curvature from the present calculation with a modified Becke-Johnson (mBJ) functional, and the squares are from a calculation by Tsirkin et al. [18] with the HSE06 hybrid functional. The bottom panel shows the quantum metric from the present calculation; the longitudinal component is represented by the solid line, and the transverse one by the dashed line. The left and bottom scales are in atomic units, while the top and right scales are in SI units.
@@
The Berry curvature near the VBM is shown in the top panel of Fig. 9 along the H–K line, as a function of the distance from the H point. Its sign depends on the handedness of the structure, and which of the two translationally-inequivalent H points is chosen. The solid line is the present calculation, and the squares are the results from a previous calculation with the HSE06 functional [18]. While the curves are similar, the values calculated with the present method are slightly smaller. We speculate the difference could be due to the use of different families of functionals: meta-GGA in the present case, hybrid functional in the previous calculation. The curvature vanishes at H by symmetry.
@@
The magnetic moments near the VBM are shown in the bottom panel of Fig. 10. At the H point the upper valence band is nondegenerate, and both moments are zero by symmetry, just like the Berry curvature in Fig. 9. As the distance from H increases along the H–K line, the spin magnetic moment saturates to the Bohr magneton, while the orbital magnetic moment, which has the opposite sign, first increases and then decreases. In both cases the variations with wavevector are on the scale of the “camel back” feature in the band energies. The same quantities are shown in the top panel for the two bands near the conduction band minimum (CBM). Those bands are degenerate at H, where they have equal and opposite nearly saturated spin magnetizations. The variation of the spin magnetization with $k$ vector is negligible in the range of the figure. Instead, the orbital magnetic moment strongly diverges near H, as the two-band model from the previous section would predict in the limit of a vanishing gap. Near H the orbital moment of the lowest band has the opposite sign of the spin polarization, but away from H it decays and changes sign. In the second lowest conduction band, the orbital moment decays and changes sign much faster. For the upper valence band, the results are similar and consistent with a previous calculation [18].
The signs of $\Omega_{z}$ and ${\mathfrak{m}}_{z}$ change depending on the handedness of the crystal, on which inequivalent H point one considers, and, in Fig. 9, on whether $q_{z}$ is on the $+z$ or $-z$ direction. The relative signs of the orbital and spin magnetic moments and Berry curvature between the different bands do not change, there is only a global sign flip. As mentioned previously, the plotted results are for the left-handed structure. With the usual $32\pi/$ 3 angle between the in-plane primitive lattice vectors, and with the first Te atom at $(u=0.269,0,0)$ in lattice coordinates, the H point of the figure is at $(1/3,1/3,1/2)$ in reciprocal lattice coordinates, at a $60^{\circ}$ angle in the basal plane with respect to the projection of the atomic positions. The $q_{z}$ is in the negative direction as indicated in the inset of Fig. 8, ${\bf k}={\bf b}_{1}/3+{\bf b}_{2}/3+(1/2-q_{z}){\bf b}_{3}$, and $\Omega_{z}$ along that line has negative values.
@@
The components of the inverse effective mass of the upper valence band of Te are shown in Fig. 11 along the H–K line. The longitudinal mass at H is electron-like while the transverse is hole-like, as expected from a saddle point.
Along that line, the longitudinal effective mass changes sign going over the “hump” of the “camel back”, and at the top of the valence band it is hole-like. The maximum is $0.0185\,\,\text{\AA}^{-1}$ away from H, where the longitudinal mass is $0.311\,\,m_{\text{e}}$ and the transverse mass is $0.129\,\,m_{\text{e}}$. The corresponding values from the LAPW mBJ calculation are $0.251\,\,m_{\text{e}}$ and $0.098\,\,m_{\text{e}}$ [16], while the experimental values are $0.220\,\,m_{\text{e}}$ and $0.108\,\,m_{\text{e}}$ [58].

### v2

@@ 3.2 Gallium arsenide
The GaAs bands near $\Gamma$ illustrate how the degenerate perturbation theory described in the previous section works in practice. In the $[111]$ direction ($\Gamma$-L), diagonalization of the $4\times 4$ matrix $A^{(1)}(\hat{\bf q}=[111],n=\text{VBM},{\bf k}=\Gamma)$ of Eq. (21) for the four-fold degenerate valence band maximum (VBM) gives two nonzero and nondegenerate band velocities with opposite signs, and a twice-degenerate vanishing band velocity (see left inset of Fig. 3). For the two nondegenerate band velocities, the masses are obtained from the relevant diagonal elements of $\tilde{A}^{(2)}([111],\text{VBM},\Gamma)$, which results from applying to $A^{(2)}([111],\text{VBM},\Gamma)$ the unitary transformation that diagonalizes $A^{(1)}([111],\text{VBM},\Gamma)$; see Eqs. (33) to (36). These two diagonal elements are identical, implying that the corresponding masses are the same. For the doubly-degenerate level with vanishing band velocity, the masses are obtained by diagonalizing the corresponding $2\times 2$ submatrix of $\tilde{A}^{(2)}([111],\text{VBM},\Gamma)$, yielding two identical masses. In the $[100]$ direction ($\Gamma$-X), $A^{(1)}([100],\text{VBM},\Gamma)$ has two doubly-degenerate nonzero band velocities with opposite signs (see middle inset of Fig. 3). For each sign, the corresponding $2\times 2$ submatrix of $\tilde{A}^{(2)}([100],\text{VBM},\Gamma)$ must be diagonalized; the final result is four identical masses as discussed previously in connection with Fig 4. Finally, in the $[110]$ direction ($\Gamma$-K), $A^{(1)}([110],\text{VBM},\Gamma)$ has four distinct band velocities: two with large absolute values, and two with small absolute values (see right inset of Fig. 3). All the masses are obtained from the corresponding diagonal elements of $\tilde{A}^{(2)}([110],\text{VBM},\Gamma)$. The fact that at $\Gamma$ the effective mass degeneracies are multiples of two is a consequence of time-reversal symmetry.
3.3 Comparison between perturbation theory and finite differences
Since effective masses can be calculated by finite differences or by perturbation theory, the advantages and disadvantages of each approach should be compared. Finite differences can be used with any code that is able to produce a band structure. Perturbation theory, on the other hand, needs an implementation like the one described here and in previous works [10, 13]. However, as can be seen from Tables 1 and 3, the best parameters for the finite differences are not obvious. If one inspects Fig. 4, it is clear that using a spacing between sampling points of $10^{-5}\,\,\text{\AA }$ will yield a very different result from a spacing of $10^{-3}\,\,\text{\AA }$. Furthermore, when one has a situation like the one depicted in the insets of Fig. 3, and only a few sampling points are used, it is not trivial how to “connect the dots” in the correct way. Perturbation theory does not need any such guesswork.
The computationally expensive step in perturbation theory is solving the Sternheimer equation. When an iterative algorithm is used, the [time-consuming] operation is applying the Hamiltonian to a trial solution. This is the same operation as in an iterative diagonalization of the Hamiltonian to obtain the eigenstates. As the number of iterative steps is of the same order of magnitude, perturbation theory has a limited overhead and, more importantly, it scales with system size in the same way as the iterative diagonalization. Finite differences has the same scaling with system size, but as it needs several Hamiltonian diagonalizations for the k-point sampling, it is computationally more expensive.
@@ 4 Gapped graphene
Gapped graphene is a system where the inversion symmetry of the two-dimensional pristine structure is broken by an external potential, typically associated with a substrate on which graphene is grown epitaxially [45, 46]. The equivalence between the two atomic sites in the primitive unit cell is broken, and as a result a gap opens in the Dirac cones at the Fermi level. The simplest tight-binding (TB) model [47, 48], with just one $p_{z}$ orbital per site, leads to a two-band low-energy model with just two parameters: the magnitude $\Delta$ of the gap and the Fermi velocity $v_{\text{F}}=\sqrt{3}at/2$ [47], with $a$ the lattice constant and $t$ the hopping integral of the TB model. Quantum-geometric quantities have simple analytical expressions in two-band models [49, 50], and the low-energy model for gapped graphene has been widely used in recent times to investigate such quantities [51].
@@
The self-consistent potential of broken-symmetry graphene was calculated using a modified Becke-Johnson meta-GGA functional [44] in a supercell configuration. The Tran-Blaha constant [44] was fixed at $c_{\text{TB}}=1.04$, since the original recipe is inadequate for isolated 2D materials. While the LDA approximation is known to underestimate the bonding energy of $\sigma$ electrons with respect to $\pi$ electrons [52], we find that the modified Becke-Johnson meta-GGA functional gives results closer to experiment. A Troullier-Martins [39, 40] pseudopotential was constructed from a local-density ground-state configuration, with a core radius of 1.3 atomic units for the $s$, $p$ and $d$ channels. The local potential was a smoothed maximum of all channels, and the nonlocal part was converted to the Kleynman-Bylander form. The self-consistent calculation used a kinetic energy cutoff of 50 Ha for the plane-wave expansion, and a $6\times 6\times 2$ uniform grid for BZ integration. As it is a slab calculation, a $6\times 6\times 1$ grid could be used instead, but the number of irreducible $k$ points would be the same, and the convergence with slab separation distance would be slightly slower. The lattice constant in the calculations was $2.456\,\,\text{\AA }$ on the graphene plane, three times that value was used in the perpendicular supercell direction. The added inversion-breaking potential was tweaked to open a gap of 0.28 eV; this is the same value that was used in Ref. [51], and is also very close to the experimental value reported for epitaxial graphene on a SiC substrate [45].
@@
Figure 5 shows the DFT-pseudopotential energy bands as a function of wavevector distance $q_{x}$ from K towards $\Gamma$, together with the bands of the low-energy model,
$E_{\pm}(q_{x},0)=\pm\frac{\Delta}{2}\sqrt{1+(q_{x}/q_{0})^{2}},$ (38)
@@
$\Omega(q_{x},0)=\pm\frac{1}{2q_{0}^{2}}\frac{1}{\left[1+(q_{x}/q_{0})^{2}\right]^{3/2}};$ (40)
@@
with $x$ in the direction from K to $\Gamma$. At K we have $g^{xx}=g^{yy}=|\Omega|/2$ [note the factor-of-two difference in Eqs. (23) and (24)], and the results from the model with the fitted $q_{0}$ are very close to the DFT values, particularly for $q_{x}\ll q_{0}$. As in the case of the Berry curvature, the differences between the DFT values for the valence and conduction bands are so small that they are not noticeable on the scale of the figure. The dependence on $q_{x}$ is very different for the longitudinal and transverse components of the quantum metric. The model fit of the DFT results is very good for the longitudinal component, but for the transverse component there is a noticeable deviation. As the system has a third dimension, the $zz$ component of the metric has a nonzero value, $g^{zz}=0.56\,\,\text{\AA }^{2}$ at $q=0$, which is very small on the scale of Fig. 6, and depends very weakly on $q_{x}$.
@@
$\Bigl{(}\frac{m_{\text{e}}}{m^{*}}\Bigr{)}^{xx}(q_{x},0)$ $=\pm\frac{\Delta}{2q_{0}^{2}}\frac{1}{\left[1+(q_{x}/q_{0})^{2}\right]^{1/2}}$ (42)
$\Bigl{(}\frac{m_{\text{e}}}{m^{*}}\Bigr{)}^{yy}(q_{x},0)$ $=\pm\frac{\Delta}{2q_{0}^{2}}\frac{1}{\left[1+(q_{x}/q_{0})^{2}\right]^{3/2}},$
with $m_{\text{e}}$ the electron mass (one in atomic units). We have omitted from the right-hand side a term $1/a_{0}^{2}E_{\text{H}}$, with $a_{0}$ the Bohr radius and $E_{\text{H}}$ the Hartree energy, as they are equal to one in atomic units. The small deviations seen in Fig. 5 between the DFT and model energy bands get magnified in the inverse masses, which are their second derivatives. For $q_{x}=0$, there is a small difference in masses between the valence and conduction bands that is not apparent on the scale of the figure. As the average effective mass was used for the value of $q_{0}$ in the DFT fit, the agreement between the model and DFT calculations is not surprising. When moving away from K towards $\Gamma$, the longitudinal and transverse inverse masses become different, the longitudinal mass decreasing at a much faster rate.
@@
where the Bohr magneton, $\mu_{\text{B}}$, has the value $1/2$ in atomic units, and we have omitted again from the right-hand side the $1/a_{0}^{2}E_{\text{H}}$ factor. In contrast with previous figures, there is a clear difference between the orbital magnetic moments of the valence and conduction bands at K, although they both have values (in atomic units) close to half the inverse mass, as predicted by the low-energy model. The sign of the orbital magnetic moment is the same for the two bands, but, like the Berry curvature, it depends on which of the two inequivalent K points is considered, or on which atom the symmetry-breaking potential is placed. A consistency check is that the orbital moment must have the same sign as the Berry curvature of the conduction band.
@@
Figure 8: The relations in Eq. (44) between the trace and determinant of the quantum metric $g$ and the magnitude of the Berry curvature $\Omega$ are illustrated by plotting differences between those quantities as a function of the distance from K towards $\Gamma$. Notice the change of scale with respect to Fig. 6. In the top panel, the dashed-black lines and solid-green lines both pertain to the low-energy model, and are obtained by inserting slightly different sets of parameters into Eq. (41). For the quantity plotted in the bottom panel, that model predicts a null value.
@@
Next, we analyze the close connection between the real and imaginary parts of the quantum-geometric tensor, i.e., between the quantum metric and the Berry curvature. For nondegenerate states and two-dimensional materials, the quantum-geometric tensor becomes a $2\times 2$ hermitian matrix. The three invariants of such a matrix are the determinant and trace of the real-symmetric part (quantum metric), and the magnitude of the imaginary-antisymmetric part (Berry curvature). From the positive semi-definiteness of the quantum-geometric tensor, one obtains the following inequalities between the three invariants [53, 54],
$\frac{1}{2}\operatorname{tr}g\geq\sqrt{\det{g}}\geq\frac{1}{2}|\Omega|.$ (44)
@@
(The first inequality is just between the arithmetic and geometric averages of the non-negative eigenvalues of the quantum metric.) Inspection of Eqs. (40) and (41) shows that the second inequality saturates for the low-energy model (this actually happens for any two-band model [54]), while the first only saturates in the limits $q_{x}\ll q_{0}$ and $q_{x}\gg q_{0}$. In Fig. 8, the quantities $\frac{1}{2}\operatorname{tr}g-\sqrt{\det{g}}$ and $\sqrt{\det{g}}-\frac{1}{2}|\Omega|$ are plotted as a function of wavevector distance from K towards $\Gamma$. Comparing the vertical scale of this figure with that of Fig. 6 shows that the inequalities are very close to being saturated, particularly in the case of $\sqrt{\det{g}}-\frac{1}{2}|\Omega|$, where the two-band model would predict a null value.
@@
The two-band low-energy model fails to take into account the presence of other bands in the DFT calculation, which are responsible for the small nonzero values displayed in the bottom panel of Fig. 8. The presence of those other bands is also responsible for the small deviations between DFT and model results seen in Figs. 5 to 7, but there the difference is only quantitative and not qualitative. To estimate the corrections to the two-band model, note that the sum-over-states expression for $Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\alpha}\rangle$ in Eq. (14) contains an energy-difference denominator. The Berry curvature and quantum metric will therefore depend on the inverse of the square of those energy differences. The smallest among them is the gap $\Delta=0.28\,\,\text{eV}$; the next relevant level (taking mirror symmetry into account) is about 9 eV away, and therefore $9/0.28\sim 32$ times farther away. The corrections to the metric and Berry curvature should therefore be of the order of $(1/32)^{2}$, or about 0.1%. Indeed, this is the factor between the vertical scale of the bottom of Fig. 8 and that of Fig. 6. In the case of the inverse mass and orbital moment, the sum-over-states expressions depend on the inverse of the energy differences (not squared); as such, deviations from the two-band model are expected to be more pronounced, of the order of 3%. That is in fact almost the value by which the orbital magnetic moments at K in Fig. 7 deviate from their average. The difference between the absolute values of the conduction and valence effective masses is 1.2%.
@@ 5 Trigonal tellurium
The self-consistent potential for the left-handed structure, with space group P3221, was calculated with the cpw2000 code using a modified Becke-Johnson meta-GGA functional [44]. The pseudopotential was relativistic Troullier-Martins [39, 40] with a core radius of 2.6 atomic units, ground state configuration, and $s$, $p$ and $d$ channels. The local potential was a smoothed maximum of all channels, and the nonlocal part was converted to the KB form [27]. The crystal lattice constants were $a=4.44\,\,\text{\AA }$ and $c=5.91\,\,\text{\AA }$, and the internal displacement parameter was $u=0.269$ [55]. The self-consistent calculation used a kinetic energy cutoff of 18 Ha for the plane-wave expansion, and a $6\times 6\times 6$ uniform grid for BZ integration (63 $\bf k$} points, once symmetry was taken into account). The Tran-Blaha parameter [44] was fixed at $c_{\text{TB}}=1.10$. This value was chosen to give a band gap of $0.312\,\,\text{eV}$, the same as a previous calculation with the HSE06 hybrid functional [19], which is close to the $0.314\,\,\text{eV}$ of a GW calculation [56], and to the experimental value of $0.323\,\,\text{eV}$ [57].
@@
Figure 9: The energy bands of Te near the band gap are shown along the H–K line, near the H point ($q_{z}=0$). The upper valence and lower conduction bands are shown in the left and right panels, respectively. On this highly detailed scale, the “camel-back” structure of the valence band is clearly seen. The inset shows the hexagonal Brillouin zone. Therein, the short cyan segment on the H–K line highlights the considered region.
@@
The valence-band maximum (VBM) of trigonal Te, relevant for the naturally $p$-doped material, is located on the H–K line, the lateral edge of the hexagonal BZ, very close to the H points, which form the vertices of the hexagonal BZ: see inset of Fig. 9. The band dispersion displays a “camel-back” feature which is clearly seen in the left panel of Fig. 9, with a local minimum at H (a saddle point in three dimensions), and a maximum about $0.02\,\,\text{\AA }^{-1}$ close to it. In the present calculation, the depth of the minimum with respect to the maximum is 1.3 meV. The experimental value is 1.1 meV [58], the calculated value with an LAPW code and also a modified Becke-Johnson functional is 1.7 meV [17], and the calculated value with the VASP code and the HSE06 functional is 0.8 meV [19]. These are very small energies, and while the order of magnitude is consistent, there is some variation associated with different computational methods and functionals.
@@
Figure 10: The Berry curvature (top panel) and quantum metric (bottom panel) of the upper valence band of Te are shown along the H–K line as a function of distance from the H point. In the top panel the solid line is the Berry curvature from the present calculation with a modified Becke-Johnson functional, and the squares are from a calculation by Tsirkin et al. [19] with the HSE06 hybrid functional. The bottom panel shows the quantum metric from the present calculation; the longitudinal component is represented by the solid line, and the transverse one by the dashed line. The left and bottom scales are in atomic units, while the top and right scales are in SI units.
@@
The Berry curvature near the VBM is shown in the top panel of Fig. 10 along the H–K line, as a function of the distance from the H point. Its sign depends on the handedness of the structure, and which of the two translationally-inequivalent H points is chosen. The solid line is the present calculation, and the squares are the results from a previous calculation with the HSE06 functional [19]. While the curves are similar, the values calculated with the present method are slightly smaller. We speculate that the difference could be due to the use of different families of functionals: meta-GGA in the present work, hybrid functional in Ref. [19]. The curvature vanishes at H by symmetry.
@@
The orbital and spin magnetic moments near the VBM are shown in the bottom panel of Fig. 11. At the H point the upper valence band is nondegenerate, and both moments are zero by symmetry, just like the Berry curvature in Fig. 10. As the distance from H increases along the H–K line, the spin magnetic moment saturates to the Bohr magneton, while the orbital magnetic moment, which has the opposite sign, first increases and then decreases. In both cases the variations with wavevector are on the scale of the camel-back feature in the band energies. The same quantities are shown in the top panel for the two bands near the conduction band minimum (CBM). Those bands are degenerate at H, where they have equal and opposite nearly-saturated spin magnetizations. The variation of the spin magnetization with $k$ vector is negligible in the range of the figure. Instead, the orbital magnetic moment strongly diverges near H, as the two-band model from the previous section would predict in the limit of a vanishing gap. Near H the orbital moment of the lowest band has the opposite sign of the spin polarization, but away from H it decays and changes sign. In the second-lowest conduction band, the orbital moment decays and changes sign much faster. For the upper valence band, the results are similar and consistent with a previous calculation [19].
The signs of $\Omega_{z}$ and ${\mathfrak{m}}_{z}$ change depending on the handedness of the crystal, on which inequivalent H point one considers, and, in Fig. 10, on whether $q_{z}$ is on the +z+z or z-z direction. The relative signs of the orbital and spin magnetic moments and Berry curvature between the different bands do not change, there is only a global sign flip. As mentioned previously, the plotted results are for the left-handed structure. With the usual $2\pi/3$ angle between the in-plane primitive lattice vectors, and with the first Te atom at $u=(0.269,0,0)$ in lattice coordinates, the H point of the figure is at $(1/3,1/3,1/2)$ in reciprocal lattice coordinates, at a $\pi/3$ angle in the basal plane with respect to the projection of the atomic positions. The $q_{z}$ is in the negative direction as indicated in the inset of Fig. 9, ${\bf k}={\bf b}_{1}/3+{\bf b}_{2}/3+(1/2-q_{z}){\bf b}_{3}$, and $\Omega_{z}$ along that line has negative values.
@@
The components of the inverse effective mass of the upper valence band of Te are shown in Fig. 12 along the H–K line. The longitudinal mass at H is electron-like while the transverse is hole-like, as expected for a saddle point. Along that line, the longitudinal effective mass changes sign going over the “hump” of the camel-back, and at the top of the valence band it is hole-like. The maximum is $0.0185\,\,\text{\AA }^{-1}$ away from H, where the longitudinal mass is $0.311\,\,m_{\text{e}}$ and the transverse mass is $0.129\,\,m_{\text{e}}$. The corresponding values from the LAPW modified Becke-Johnson calculation are $0.251\,\,m_{\text{e}}$ and $0.098\,\,m_{\text{e}}$ [17], while the experimental values are $0.220\,\,m_{\text{e}}$ and $0.108\,\,m_{\text{e}}$ [59].